OpenStudy (halpppmehhhh):

will medal and fan! 7/4x - 5/6 = 5/8(x-2/3) whats the value of x? (The answer is in a fraction or mixed fraction form)

OpenStudy (mathmale):

In your 7/4x - 5/6 = 5/8(x-2/3), 5/6 and 2/3 are clearly stand-alone fractions. But your 7/4x needs clarification: do you mean 7x/4 or do you mean 7/(4x)?

OpenStudy (halpppmehhhh):

I meant exactly what i put. 7/4 is a fraction but on my paper it shows the x right next to the fraction. I don't understand this.

OpenStudy (mathmale):

Again I am pointing out that there's ambiguity here. If 7/4 is a fraction, and is also the coefficient of x, please write this term as (7/4)x or as 7x/4.

OpenStudy (halpppmehhhh):

Ohh okay

OpenStudy (mathmale):

Let's transform your 7/4x - 5/6 = 5/8(x-2/3) into (7/4)x - 5/6 = 5/8(x-2/3). It'd be so much easier to do this problem if we could get rid of all the fractions. If you look carefully you'll see that the fractions 7/4, 5/6 and (5/8)(2/3) all have the common denominator 24. If you have some other way of removing the fractions, by all means go ahead with it.